1. A solution is prepared by mixing 2.80 mol of acetone with 1.45 mol of cyclohexane at 30°C.
Calculate the Xacetone and Xcyclohexane in the vapor above the solution.
P°acetone = 229.5 torr and P°cyclohexane = 97.6 torr.
Xacetone in the vapor =
Xcyclohexane in the vapor =
1
Expert's answer
2020-02-25T04:51:50-0500

X1=P1/(P1+P2)

Xacetone=229.5/(229.5+97.6)=0.70=70%

Xcyclohexane=97.6/(229.5+97.6)=0.3=30%